OpenStudy (superhelp101):

factor 4x^2-17x+4 by grouping

OpenStudy (superhelp101):

@vishweshshrimali5 do you think you could help meh?

OpenStudy (vishweshshrimali5):

\[\large{4x^2 - 17 x+ 4}\] \[\large{= 4x^2 - 16 x - x + 4}\] \[\large{= 4x(x - 4) - (x-4)}\] Can you solve after this ?

OpenStudy (vishweshshrimali5):

Any doubt in any step I wrote above ?

OpenStudy (tkhunny):

I so hate "factor by grouping". This method is rarely of any value whatsoever - except in a section entitled "factor by grouping". It usually helps me to solve some other way so I can see what "grouping" might be useful.

OpenStudy (superhelp101):

yep! (4x-1)(x-4). But I feel there's something wrong

OpenStudy (vishweshshrimali5):

Well nothing wrong I think :)

OpenStudy (vishweshshrimali5):

\[\large{(4x-1)(x-4) = 4x^2 - 16x - x + 4}\] \[\large{= 4x^2 - 17x +4}\] Correct !

OpenStudy (superhelp101):

You're right! :) but i am supposed to get: